OpenStudy (anonymous):

use the distributive property to find each product 27(2 1/3)

OpenStudy (anonymous):

63

OpenStudy (anonymous):

OK ur answer is following

OpenStudy (anonymous):

49/3

OpenStudy (anonymous):

HERE this is the distributive property of multiplication under addition

OpenStudy (anonymous):

2 1/3 is an mixed number

OpenStudy (anonymous):

NOW question is 27 ( 2 + 1/3) i think

OpenStudy (anonymous):

a ( b+ c) = ab + bc (27*2) + (27*1/3) = 54 + 9 = 63